OpenStudy (anonymous):

what is the sum of 3/8 and 1/4

OpenStudy (unklerhaukus):

\[3/8+1/4\]\[=3/8+2/8\]\[=5/8\]

OpenStudy (anonymous):

show your work

OpenStudy (anonymous):

that is confusing